A solution is made by dissolving 0.539 mol of nonelectrolyte solute in 775 g of benzene. Calculate the freezing point, Tf, and boiling point, Tb, of the solution.
Answer #1

Apply:

dTf = -Kf*m

m = mol solute / kg solvnet = 0.539/0.755 = 0.71390 molal

dTf = -1.86*0.71390 = -1.327854 °C

Tf = -1.327854 °C

for boiling point

dTb = Kb*m = 0.512*0.71390

Tb= 100+0.3655168 = 100.3655168 °C